OpenStudy (anonymous):

integral of cos6x dx plz help!!

OpenStudy (anonymous):

\[\frac{ \sin(6x) }{ 6 } + C\] because its derivative is cos(6x)

zepdrix (zepdrix):

:O

OpenStudy (anonymous):

how did u get that?? @Euler271

OpenStudy (anonymous):

???? @zepdrix

OpenStudy (anonymous):

what's the derivative of cos(6x)?

OpenStudy (anonymous):

-sin6x ???? @Euler271

OpenStudy (anonymous):

-6sin6x

OpenStudy (anonymous):

the integral is the inverse operation of the derivative. so the integral of -6sin6x must be cos6x. tbh i dont even know how to explain it lol. i hope this helps

OpenStudy (anonymous):

yea sorry!! I get u! thanks!@Euler271

OpenStudy (anonymous):

^_^

zepdrix (zepdrix):

That's a good method to get comfortable with, just sort of like.. advanced guessing. You know that the derivative will produce an extra factor of 6 because of the 6x inside of the trig function. So when we integrate, we need to divide by a factor of 6. If this is still confusing though, you can certainly perform a u-substitution:\[\Large \int\limits \cos6x\;dx,\qquad\qquad u=6x\]